A survey asked how many children people have. The survey results are given in the histogram above. List the raw data (survey responses).

" PLEASE EXPLAIN TO ME HOW YOU GOT THE ANSWER; THE STEPS PLEASE AND THANK YOU "

data A survey asked how many children people have. The survey results are given in the histogram above. List the raw data (survey responses)
Answer #1

Number of children =data * no of frequency = 1*2 + 2*5 + 3*3 + 4*4 = 37 (ans)